OpenStudy (cakt07):

What is 3.25% written as a fraction in lowest terms?

OpenStudy (cakt07):

@orion92898

OpenStudy (anonymous):

3.25% = 3.25/100 = 3.25*4 / (100*4) = 13/400

OpenStudy (anonymous):

13/4

OpenStudy (anonymous):

orion u are smart

OpenStudy (anonymous):

^^^ That is what you get if you take the 0 0 off my answer sorry would get in trouble if I gave a direct answer @jswagger thanks.

OpenStudy (anonymous):

welcome